# Holographic symmetry algebra for the MHV sector revisited

### v1

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@ 3 Knizhnik-Zamolodchikov type null states for negative helicity gravitons
[…]
@@
$\begin{split}-\bar{z}_{12}\bar{H}^{-3}_{\frac{5}{2},-\frac{1}{2}}G^{+}_{\Delta_{1}}\left(z_{1},\bar{z}_{1}\right)=&\bar{z}_{12}\bigg{[}L_{-1}G^{-}_{\Delta_{1}-3}(z_{1},\bar{z}_{1})+H^{0}_{0,-1}H^{1}_{-\frac{3}{2},\frac{1}{2}}G^{-}_{\Delta_{1}-4}(z_{1},\bar{z}_{1})\\ &+H^{0}_{-1,0}G^{-}_{\Delta_{1}-3}(z_{1},\bar{z}_{1})+\Delta_{1}~H^{1}_{-\frac{3}{2},-\frac{1}{2}}G^{-}_{\Delta_{1}-4}(z_{1},\bar{z}_{1})\bigg{]}\\ &\end{split}$ (17)
@@ Appendix A Algebra for graviton
$\begin{split}\bar{H}^{k}(z,\bar{z})=\lim_{\Delta\rightarrow k}(\Delta-k)G^{-}_{\Delta}(z,\bar{z}),~~~k=-3,-4,-5,...\end{split}$ (36)
@@ Appendix C The $KZ$ type null state for gluon
Multiplying with $-if^{abd}$ and using $f^{aa_{1}b}f^{aa_{1}c}=C_{A}\delta^{bc}$, we will get the following relations

### v3

@@ Abstract
We revisit the holographic symmetry algebra in the MHV sector. We find an infinite dimensional Abelian symmetry algebra whose generators are the conformally soft negative helicity gravitons and gluons. So the complete symmetry algebra in the MHV graviton sector is a semideirect product of the $w_{1+\infty}$ algebra and the infinite dimensional Abelian algebra. Similarly in the MHV gluon sector the symmetry algebra is a semidirect product of the $S$ algebra and the infinite dimensional Abelian algebra. The extended symmetry algebra has some use. For example, it is known for sometime that an $n$ point MHV amplitude satisfies (n2)(n-2) Knizhnik-Zamolodchikov (KZ) type equations. So two equations are missing. We show that the extended symmetry algebra has additional null states whose decoupling give rise to the two missing equations.
@@ 1 Introduction
Symmetry plays an important role in physics and identifying all the non-trivial symmetries of the S matrix in asymptotically flat spacetime(AFS) still remains an open problem. Celestial Holography gerger:2017zoo ; Strominger:2013lka ; Banerjee:2020zlg ; Banerjee:2021cly ; Guevara:2021abz ; Strominger:2021lvk ; Ball:2021tmb ; Adamo:2021lrv ; Guevara:2025tsm ; Banerjee:2023zip ; Banerjee:2020vnt ; Banerjee:2023bni is a very useful framework to address such issues. In this paper we revisit the soft symmetries or the Holographic Symmetries Algebra Guevara:2021abz ; Strominger:2021lvk for the (tree level) MHV scattering amplitudes. In Banerjee:2020zlg , it was shown that the MHV graviton scattering amplitudes are completely determined by the symmetry algebra which is a semidirect product of a $SL_{2}(\mathbb{R})$ current algebra and supertranslations. It works in the following way. One can show that the representation of the symmetry algebra, which is relevant for the MHV amplitudes, has null states. These null states are similar to the Knizhnik-Zamolodchikov (KZ) null states that are there in the WZW models. The existence of these null states lead to differential equations for the Celestial correlators which one can solve to get the MHV amplitudes. However, there is a puzzle. For an $n$-point MHV amplitude one gets only (n2)(n-2) KZ type differential equations. This is unlike the case of the WZW model where there are $n$ KZ equations for an $n$ point correlation function. This puzzle is not solved by the extended symmetry algebra Strominger:2021lvk of the MHV sector which is $w_{1+\infty}$.
@@ 3 Knizhnik-Zamolodchikov type null states for negative helicity gravitons
The symmetry algebra (14) admits null states. In this section we are interested in the Knizhnik-Zamolodchikov (KZ) type null states which contain the $L_{-1}$ descendent of the (hard) graviton operator on the celestial sphere. Such KZ type null states in the MHV sector were found in Banerjee:2020zlg . However, if we use only the symmetry algebra generated by the positive helicity (conformally soft) gravitons then we get only (n2)(n-2) KZ type equations for an $n$-point MHV amplitude. The (n2)(n-2) null states contain the $L_{-1}$ descendant of the positive helicity hard graviotn only. So the list of KZ type null states obtained in Banerjee:2020zlg was incomplete. In this section we fill in this gap and find the two missing null states using the infinite Abelian symmetry algebra generated by the conformally soft negative helicity gravitons in the MHV sector.
@@
[…]
@@ Appendix A Algebra for graviton
[…]
@@ Appendix C The $KZ$ type null state for gluon
Multiplying with $if^{abd}$ and using $f^{aa_{1}b}f^{aa_{1}c}=C_{A}\delta^{bc}$, we will get the following relations